OpenStudy (anonymous):

challenge question find \(\sum _{i=1}^{(40)}i^7 \mod 100 \) i wont be here for at least 4 hours so have fun :P

ganeshie8 (ganeshie8):

0 ?

ganeshie8 (ganeshie8):

not really a challenge :P

OpenStudy (anonymous):

haha m but why zero :P

OpenStudy (anonymous):

i was gonna say 99 instead of 7 but though giving the chance for ppl how likes to solve it without back ground

ganeshie8 (ganeshie8):

can we say \[\begin{array}{} \sum\limits_{i=1}^{\phi(a)} i^n \equiv 0 \pmod a &:& \phi(a) \not |~ n \\~\\ \sum\limits_{i=1}^{\phi(a)} i^n \equiv -1 \pmod a &:& \phi(a) | n \\~\\ \end{array} \]

OpenStudy (anonymous):

ermm na see for n>101 this does not work :3

OpenStudy (anonymous):

Did someone call me ?

OpenStudy (anonymous):

:O speaks of the devil !

OpenStudy (anonymous):

Well ?

ganeshie8 (ganeshie8):

it should work for all powers let me check

OpenStudy (anonymous):

ok k tbh n>99 does not work

ganeshie8 (ganeshie8):

why ?

OpenStudy (anonymous):

:D can u read its a challenge

ganeshie8 (ganeshie8):

it works for all powers when the moduli is prime, right ?

OpenStudy (anonymous):

:P yes when the prime <100 and works for other integers however with phi its not true u need to find some other easy way

OpenStudy (mathmath333):

100 is not prime , i was about to apply the the conclusion from last question lol

OpenStudy (anonymous):

k have to go , have fun !

ganeshie8 (ganeshie8):

ok you too have fun :)

OpenStudy (anonymous):

playing with snow :3

ganeshie8 (ganeshie8):

yeah 100 is not a prime but we can still use the previous result by considering \(\phi (100)\)

ganeshie8 (ganeshie8):

@Marki do you have a proof or something for why it doesn't work for composite numbers ?

ganeshie8 (ganeshie8):

Oh I see. composite integers need not have a primitive root and the previous proof breaks

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

i still dont get why after 4 they are divisible by 100

OpenStudy (anonymous):

@ganeshie8 try lol

OpenStudy (mathmath333):

oh wait i think its faulty lol

OpenStudy (mathmath333):

is their a partial sum formula to that series

OpenStudy (anonymous):

i dnt know all formulas that solves it but u can try why not

OpenStudy (kainui):

Here's one way of doing it that is no doubt inferior to what ganeshie and markie have in mind... But it's fun so I'll start you all out: \[\Large \sum_{i=1}^{40}i^8-(i-1)^8=40^8 \] Take the telescoping series a power above, then expand the binomial **scribbles half of pascal's triangle** \[\sum_{i=1}^{40}i^8-(i^8-8i^7+28i^6-56i^5+70i^4-56i^3+28i^2-8i+1)=40^8\] change some junk around \[\sum_{i=1}^{40}8i^7-28i^6+56i^5-70i^4+56i^3-28i^2+8i-1 =40^8\] \[\sum_{i=1}^{40}i^7 =\frac{1}{8}40^8 + \frac{1}{8}\sum_{i=1}^{40}28i^6-56i^5+70i^4-56i^3+28i^2-8i+1\] Yay we decreased the order of the sum by 1, let's just repeat this a few more times...

ganeshie8 (ganeshie8):

i think it works if we run the sum over numbers that are relatively prime to 100

OpenStudy (dan815):

hahaha what is with u and telescoping series now

ganeshie8 (ganeshie8):

Kai's method works always no matter what ! evaluate the sum and find the remainder. pretty straightforward :)

OpenStudy (anonymous):

haha yeah indeed

OpenStudy (anonymous):

that's why i chose 7 not 99 :P

ganeshie8 (ganeshie8):

@Marki I see the remainder of the sum cannot be 0 or -1 for all powers because fermat's little thm requires the base to be relatively prime to the moduli and there will be many integers between 1 and 40 that are not relatively prime to 100

OpenStudy (dan815):

0 for all odd powers of 7 and =/=0 for all even powers of 7?

OpenStudy (anonymous):

ok here is hint \(\sum _{i=1}^{(40)}i^7 =0 \mod 5\) \(\sum _{i=1}^{(40)}i^7 =0 \mod 2\)

ganeshie8 (ganeshie8):

but if you run the sum over integers that are relatively prime to 100, the remainder will be always 0 or -1 for all powers : \[\sum \limits_{\gcd(i,100) = 1} i^{n} \pmod {100}\]

ganeshie8 (ganeshie8):

thats clever @Marki xD \(5 \not |~ 7\) and \(2 \not | ~ 7\) so both sums evaluate to 0

ganeshie8 (ganeshie8):

how are you going to get to mod 100 from them ?

OpenStudy (anonymous):

since gcd(2,5)=1 then \(\sum _{i=1}^{(40)}i^7 =0 \mod 10\)

ganeshie8 (ganeshie8):

Yes

OpenStudy (anonymous):

\(\sum _{i=1}^{(40)}i^7 =10k\)

ganeshie8 (ganeshie8):

yes thats the dead end.

OpenStudy (anonymous):

:D why dead we need to show k=10m :P

ganeshie8 (ganeshie8):

idk how ?

OpenStudy (anonymous):

actually i did it with primitive roots as u did in last question

OpenStudy (kainui):

AHA! We can represent all the i's in terms of a basis: \[\Large i= a*10^1+b*10^0\] Raise this to the 7th power mod 100 we have: \[(a*10+b)^7 \mod 100 \equiv 70 ab^6 + b^7\]

OpenStudy (anonymous):

lol i got stuck with my own solution :'(

OpenStudy (kainui):

We can already remove every multiple of 10 from the sum leaving us with 1 to 9, 11 to 19, etc... and representing it this way we really have: \[\sum_{i=1}^{40} 70ab^6+b^7= 70 \sum_{a=1}^3 \sum_{b=1}^9 ab^6 + 4 \sum_{b=1}^9 b^7\] This is feeling matrixy but the idea seems fairly simple, b is the first digit, and a is the second digit.

OpenStudy (kainui):

Expressed this way, it's obvious now that when a=2 and b=5 that we can drop out those terms from the first "double summation" since they will just be 70*2*5=700 which is 0 mod 100. \[\ 70 \sum_{b=1}^9( 1b^6+3b^6) + 4 \sum_{b=1}^9 b^7\] \[\ 280 \sum_{b=1}^9b^6 + 4 \sum_{b=1}^9 b^7\] \[\ 280 \left( \sum_{b=1}^4b^6+ \sum_{b=6}^9b^6 \right) + 4 \sum_{b=1}^9 b^7\] Then remembering this is mod 100, change the 280 into 80. \[\ 80 \left( \sum_{b=1}^4b^6+ \sum_{b=6}^9b^6 \right) + 4 \sum_{b=1}^9 b^7\] Actually I also noticed just now that when b is 5 it will give us 100 on the other b^7 summation too, so we can remove it there as well: \[\ 80 \left( \sum_{b=1}^4b^6+ \sum_{b=6}^9b^6 \right) + 4 \left( \sum_{b=1}^4b^7+ \sum_{b=6}^9b^7 \right)\]

OpenStudy (kainui):

We can recombine and factor around to get both summations like this: \[\Large 4 \sum_{b=1 \ or \ b=6}^{4 \ or \ 9}b^6(20+b) \] Which, let's be honest, is simple to calculate because it's only 8 terms.

OpenStudy (anonymous):

:|

OpenStudy (anonymous):

i was like gonna say there is no order 7 in modulo 100

OpenStudy (anonymous):

@ganeshie8 right ? so this equivalent work \(r^{0.7}+r^{1.7} + r^{2.7} + \cdots + r^{7(\phi(100)-1)} = \dfrac{1-r^{(7(\phi(100) ) )}}{1-r^7} \equiv 0 \pmod {100}\) but now not sure

ganeshie8 (ganeshie8):

Here is an insanely short solution : \(\sum\limits_{i=1}^{40} i^7 \equiv \sum\limits_{i=1}^{40} i^3+5k\equiv \sum\limits_{i=1}^{40} i^3 \equiv 0 \pmod{5^2} \) \(\sum\limits_{i=1}^{40} i^7 \equiv \sum\limits_{i=1}^{40} i \equiv 0 \pmod{2^2} \) combining them gives \(\sum\limits_{i=1}^{40} i^7 \equiv 0 \pmod {5^22^2}\) XD

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

\[r^{0.7}+r^{1.7} + r^{2.7} + \cdots + r^{7(\phi(100)-1)} = \dfrac{1-r^{(7(\phi(100) ) )}}{1-r^7} \equiv 0 \pmod {100}\] this is tricky business : how do you know a primitive root exists for 100 ? even if it exists, the bases in question {1,2,...,40} are not all relatively prime to 100. so your logic wont work.

OpenStudy (anonymous):

cc ?

OpenStudy (anonymous):

well hmm xD yeah i saw its wrong lately

ganeshie8 (ganeshie8):

did u get my previous solution, im in love wid fermat already xD

OpenStudy (anonymous):

yes it was nice !! first i thought like u \(\sum _{i=1}^{\phi (n)} i^p \mod n ~~~~ \forall p<100\)

OpenStudy (anonymous):

is zero , but then i dint know hw t pr0ve :|

ganeshie8 (ganeshie8):

it will be 0 or -1 if all the i's are relatively prime to 100

OpenStudy (anonymous):

naa forget about -1 i found counter example xD

ganeshie8 (ganeshie8):

when ? it is an identity and we have proved it in your previous question

OpenStudy (anonymous):

naa we proved Fermat sum ,not Euler sum

ganeshie8 (ganeshie8):

both are same all you need to do is let the i's be relatively prime to 100

ganeshie8 (ganeshie8):

Below should be true whenever a primitive root exists for \(a\) : \[ \begin{array}{} \sum\limits_{\gcd(i,a)=1} i^n \equiv 0 \pmod a &:& \phi(a) \not |~ n \\~\\ \sum\limits_{\gcd(i,a)=1} i^n \equiv \phi(a) \pmod a &:& \phi(a) | n \\~\\ \end{array} \]

ganeshie8 (ganeshie8):

notice when \(a\) is prime, you get the theorem that we proved in your previous question

ganeshie8 (ganeshie8):

let me know if u want a proof or examples

OpenStudy (anonymous):

second part works but im nt convinced about the first one

ganeshie8 (ganeshie8):

gcd(i, 100) must be 1 you're running over first 40 positive integers, so obviously it wont work

OpenStudy (anonymous):

maybe its like this \(\begin{array}{} \sum\limits_{\gcd(i,a)=1} i^n \equiv 0 \pmod a &:& \phi(a) \not |~ n \\~\\ \sum\limits_{\gcd(i,a)=1} i^n \equiv \phi(n) \pmod a &:& \phi(a) | n \\~\\ \end{array}\)

OpenStudy (anonymous):

ok i'll work ltr on it , cya

ganeshie8 (ganeshie8):

Okay, il give you a chocolate if you find a counter example for above ^ :)

ganeshie8 (ganeshie8):

\[\ 80 \left( \sum_{b=1}^4b^6+ \sum_{b=6}^9b^6 \right) + 4 \left( \sum_{b=1}^4b^7+ \sum_{b=6}^9b^7 \right)\] \[\sum_{b=1}^4 \left( 80b^6 + 4b^7 + 80(b+5)^6 + 4(b+5)^7\right)\] @Kainui just 4 terms now :) and it does give correct answer after playing so much https://www.wolframalpha.com/input/?i=%5Csum%5Climits_%7Bb%3D1%7D%5E4+%2880b%5E6%2B4b%5E7+%2B+80%28b%2B5%29%5E6+%2B+4%28b%2B5%29%5E7%29+mod+100 xD

OpenStudy (anonymous):

chocolate idea excited me but u sound so sure (u dont seems like giving it ) :| so i would believe u and say ok no counter example but u give me one true statement about it ok ?

OpenStudy (anonymous):

why u add this "^" when u end a statement ? what does it even mean ?

OpenStudy (anonymous):

@Kainui approach is amazing though

OpenStudy (kainui):

Back, my internet died right after I posted my answer. there's too much for me to read now, maybe when I wake up I'll read more lol.

OpenStudy (anonymous):

@ganeshie8 which case work ?

OpenStudy (kainui):

Oh ganesh has a good idea! \[\Large 4 \sum_{b=1}^{4}b^6(20+b) + (b+5)^6(25+b) \] And yet more terms are cut out after distributing if you recognize the 25 and 4, so we have: \[\Large 4 \sum_{b=1}^{4}b^6(20+b) + b (b+5)^6 \] Expanding the binomial b+5 we see that the 5^2 terms will disappear as well! Getting warmer! \[\Large 4 \sum_{b=1}^{4}b^6(20+b) + b (b^6+6*5b^5)\] Redistribute and we have: \[\Large 4 \sum_{b=1}^{4}50b^6+2b^7\] And now we are left with just a few terms after seeing 4*50=0 mod 100..! (hope it's right XD) \[\Large 8 \sum_{b=1}^{4}b^7 = 8(1+2^7+3^7+4^7) \] But I still haven't reached the answer with this ridiculous wandering.Hmmm... Powers of 2 lol.

OpenStudy (anonymous):

lol its like watching good match :P

ganeshie8 (ganeshie8):

like watching a good test match between australia and england :)

OpenStudy (anonymous):

naa it would be boring :3 i mean real match

OpenStudy (kainui):

Hah! I'll take three of the terms and reverse their powers around into a mini geometric series! XD \[\large 1+2^7+4^7 = (2^7)^0+(2^7)^1+(2^7)^2 = \frac{(2^7)^3-1}{(2^7)-1}\] (we really should write this in base 2 shouldn't we I'll just look at the numerator since we? XD) Here I just wrote the numerator since we know this is a whole number from the left that means this has a factor of 2^7-1 \[\large 2^{21}-1=111111111111111111111_2\] Let's keep playing, I just thought I should share before I keep going wherever we end up @_@

OpenStudy (kainui):

So yeah on with it in base 2. XD \[\frac{111111111111111111111}{1111111}= \frac{1111111*100000010000001}{1111111} = \] \[\LARGE 100000010000001 = nevermind...\]

ganeshie8 (ganeshie8):

repunits are pretty neat stuff 2^7 = 28 so 1+2^7 + 4^7 = 1 + 28(1+28)

ganeshie8 (ganeshie8):

@ganeshie8 which case work ? what are you refering to @Marki

OpenStudy (anonymous):

nvm